OpenStudy (tanabugg):

A survey about the student government program at a school finds the following results: 110 students like the program. 120 students think the program is unnecessary. 210 students plan on running for student government next year. If a circle graph were made from this data, what would the measure of the central angle be for the group that LIKES the program? 25 degrees 90 degrees 98 degrees 172 degrees

OpenStudy (alexishope47):

Using this equation, find x: Children who like it/total children=measure of angle/total degrees in circle 110/440=x/360 This should give you what you are looking for. I took geometry last year and, correct me if I'm wrong but, a central angle has the same degree as the arc it creates so find x and that should be your answer.

OpenStudy (alexishope47):

@tanabugg let me know if you are still confused.

OpenStudy (tanabugg):

So 25% @alexishope47 thanks so much!

OpenStudy (alexishope47):

How did you solve for x?

OpenStudy (alexishope47):

I got a different answer.

OpenStudy (tanabugg):

I did 110/ 440= 0.25.

OpenStudy (alexishope47):

Try solving by cross multiplying. Like they are fractions. 110/440=x/360 (360)*(110)=440x That's how I always did it in school.